OpenStudy (anonymous):

The population of a town was 47,000 in 1950 and 83,000 in 2010. Assume the rate of population increase is constant. What will the population be in 2015? a.85,400 b.89,000 c.83,600 d.86,000

OpenStudy (ranga):

2010 83,000 1950 47,000 Difference: 60 years 36,000 In 60 years, the population increased by 36,000 In 5 years (that is, from 2010 to 2015) how much will the population increase? Use proportions to figure out the increase in population and add it to 83,000 to find the population in 2015.

OpenStudy (anonymous):

that did not help

OpenStudy (ranga):

Let x be the increase. Use proportions: 36,000 in 60 years is in proportion to x in 5 years. 36,000 / 60 = x / 5 Solve for x. Add it to 83,000.

OpenStudy (anonymous):

im in 8th grade so plz explain it so i understand it

OpenStudy (ranga):

In 60 years the population increases by 36,000 Therefore, in 1 year, the population will increase by 36,000 / 60 = 600 In 5 years the population will increase by 600 * 5 = 3,000 So the population of 83,000 in the year 2010 will increase to 83,000 + 3,000 in 2015. 83,000 + 3,000 = 86,000.